# Bounds on the number of squares in recurrence sequences

### v1

@@ 3.4. Miscellaneous Lemmas
$\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}>\frac{21.12\sqrt{d}\,\left(\left|N_{\alpha}\right|/4\right)}{8\sqrt{\left|N_{\alpha}\right|}}\geq 0.66\sqrt{d},$
where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.35) to establish the first inequality. For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
@@
For each pair, we check (3.36) and (3.37) by computing their left-hand sides for all $k\neq 0$ such that $20.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<$ 2 and $30021.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<$ 300. No counterexamples to (3.36) and (3.37) were found. The script took 112 seconds to run on a laptop with an Intel i7-9750H processor and 16gb of RAM.
@@
Our choice of the lower bound for $d$ comes from an example of $E<1$ for $d=104$: $a=9$, $d=104$, $N_{\alpha}=-23$, $k=-1$, $x_{k}=-61$ and $y_{k}=11$ where $E=0.973\ldots$. ∎
@@
Conjecture 1.2 would follow for all sequences with $b=1$ and $-N_{\alpha}$ a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$. This is due to Lemma 3.5.
@@ 4.1. Prerequisites
We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$.
@@
This arises from Lemma 3.12 and will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof, in Subsection 4.6.
@@
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.06.$
Thus we can apply Lemma 3.6(a) with $c_{1}=0.06$ to find that
(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
@@
(4.5) $E$ $=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}>\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{4e^{1.68}\left|N_{\alpha}\right|}$
@@
From Lemma 3.5(c), $|g|\mathcal{N}_{d^{\prime},4}\geq 2$ and (4.1), we have $E>1$, as required for its use with Lemma 2.1.
Similarly, using (4.3) and (3.37) in Lemma 3.12 we have
@@
Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.8(a), we can take
(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.46\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
@@ 4.2. $r_{0}=1$ and $\zeta_{4}p/q\neq p_{1}/q_{1}$ for all $4$-th roots of unity, $\zeta_{4}$
From equation (4.2), along with Lemma 2.1(b) and (3.25) with $b=1$, we have
(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.99\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.99(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{3.99(1-c)}{0.89\left(21.47\sqrt{d}\,y_{k}/\left(|g|\mathcal{N}_{d^{\prime},4}\right)\right)^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
(4.11) $\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}>y_{\ell}^{3}.$
@@
(4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(N_{\alpha}f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
@@
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.7(a), we find that
$57.6^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
@@
By (4.1), we have $0.250.24/\left((1-c)d^{1/2}\right)<$ 0.25 if $c<0.9$. We will see in Subsection 4.4 that $c=0.75$ is a good choice. So $y_{k}<\left|N_{\alpha}\right|$/4, which contradicts Lemma 3.5(c). Hence this case cannot hold.
@@ 4.3. $r_{0}=1$ and $\zeta_{4}p/q=p_{1}/q_{1}$ for some $4$-th root of unity, $\zeta_{4}$
$=\left|\frac{N_{d,4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d,4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t}}{2g}\right)^{r_{0}}\right|$
$=\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
(4.13) $\geq\left|\frac{1}{Y_{1,4,r_{0}}(\omega)}\right|\left|(\omega-1)^{2r_{0}+1}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
@@
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.291}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
@@
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{1.16}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1},$
@@
(4.15) $1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>y_{\ell}.$
@@
$6.92\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.6\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.6\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible.
@@ 4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
$\left(y_{k}y_{\ell}\right)^{1/4}$ $>\frac{0.995c\left|x_{k}\right|}{0.46\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}$
$>\frac{2.1c\sqrt{d}\,y_{k}}{\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}.$
@@
(4.17) $y_{\ell}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}.$
@@
$>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{12}c^{12}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{12r_{0}}d^{6r_{0}}y_{k}^{12r_{0}-3}.$
@@
(4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
@@
$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
@@
$\frac{0.00079\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$

### v2

@@ 3.4. Miscellaneous Lemmas
$\frac{21.12\sqrt{d}\,y_{k}}{|g|\mathcal{N}_{d^{\prime},4}}\geq\frac{21.12\sqrt{d}\,\left(\left|N_{\alpha}\right|/4\right)}{8\sqrt{\left|N_{\alpha}\right|}}\geq 0.66\sqrt{d},$
where we use $|g|\mathcal{N}_{d^{\prime},4}\leq 8\sqrt{\left|N_{\alpha}\right|}$ from (3.36) and Lemma 3.5(c) with $u\geq 1$ to establish the first inequality. For $d\geq 109,000$, we find that $2170.66\sqrt{d}>$ 217 holds.
@@
For each pair, we check (3.37) and (3.38) by computing their left-hand sides for all $k\neq 0$ such that $20.1832\cdot 2^{1+\min\left(v_{2}(b)/2,2\right)}\sqrt{d}y_{k}/b<$ 2 and $30021.12y_{k}/2^{1+\min\left(v_{2}(b)/2,2\right)}<$ 300. No counterexamples to (3.37) and (3.38) were found. A PARI/GP program took 112 seconds to run on a Windows 10 laptop with an Intel i7-9750H processor and 16GB of RAM.
@@
Conjecture 1.2 would immediately follow when $-N_{\alpha}$ is a square, if we could replace the lower bound for $y_{k}$ in Proposition 4.1 with $\max\left(1,\left|N_{\alpha}\right|/4\right)$. This is due to Lemma 3.5.
@@ 4.1. Prerequisites
We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$. So $y_{\ell}>y_{k}\geq 4\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right\}$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
@@
This is the condition in Lemma 3.13, which will allow us to bound $E$ and $Q$ from below. This assumption shall be removed at the end of the proof in Subsection 4.6.
@@
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
Thus we can apply Lemma 3.7(a) with $c_{1}=0.127$ to find that
(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
@@
(4.5) $E$ $=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|u_{1}\right|+\sqrt{u_{1}^{2}-t^{\prime}u_{2}^{2}}\right|}{\mathcal{D}_{4}u_{2}^{2}|t^{\prime}|}=\frac{|g|\mathcal{N}_{d^{\prime},4}\left|\left|2x_{k}\right|+2\sqrt{x_{k}^{2}-N_{\alpha}}\right|}{4e^{1.68}\left|N_{\alpha}\right|}$
@@
From (3.38) in Lemma 3.13, we have $E>1.13>1$, as required for its use with Lemma 2.1.
Similarly, using (4.3) and (3.38) in Lemma 3.13, we have
@@
Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
@@ 4.2. $r_{0}=1$ and $\zeta_{4}p/q\neq p_{1}/q_{1}$ for all $4$-th roots of unity, $\zeta_{4}$
From the equality in (4.2), along with Lemma 2.1(b) and (4.9), we have
(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{3.959(1-c)}{0.89\left(21.47\sqrt{d}\,y_{k}/\left(|g|\mathcal{N}_{d^{\prime},4}\right)\right)^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
(4.11) $\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}\left(\frac{0.45}{1-c}\right)^{4}\left(\frac{21.47}{|g|\mathcal{N}_{d^{\prime},4}}\right)^{4r_{0}}d^{2r_{0}-2}y_{k}^{4r_{0}+1}>y_{\ell}^{3}.$
@@
(4.12) $y_{\ell}^{3}<545(1-c)^{-4}\left(\left|N_{\alpha}\right|f_{k}f_{\ell}\right)^{2}y_{k}^{5}.$
@@
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
$57.32^{3}\left(\frac{d}{\left|N_{\alpha}\right|}\right)^{6}y_{k}^{4}<545(1-c)^{-4}N_{\alpha}^{2},$
@@
By (4.1), we have $0.250.24/\left((1-c)d^{1/2}\right)<$ 0.25 if $c<0.9$. We will see in Subsection 4.4 that $c=0.75$ is a good choice. So $y_{k}<\left|N_{\alpha}\right|/4$, which contradicts parts (b) and (c) of Lemma 3.5. Hence this case cannot hold.
@@ 4.3. $r_{0}=1$ and $\zeta_{4}p/q=p_{1}/q_{1}$ for some $4$-th root of unity, $\zeta_{4}$
$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
$=$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
(4.13) $\geq$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
@@
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
@@
$\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>\frac{1.154}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1},$
@@
(4.15) $1.734r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>y_{\ell}.$
@@
$6.96\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3}>y_{\ell}>57.32\frac{d^{2}}{\left|N_{\alpha}\right|^{2}}y_{k}^{3}>57.32\frac{d}{\left|N_{\alpha}\right|}y_{k}^{3},$
since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
@@ 4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
$\left(y_{k}y_{\ell}\right)^{1/4}$ $>\frac{0.995c\left|x_{k}\right|}{0.458\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}$
$>\frac{2.103c\sqrt{d}\,y_{k}}{\sqrt{\left|N_{\alpha}\right|f_{k}f_{\ell}}}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}\sqrt{d}\,y_{k}}{\left|N_{\alpha}\right|}\right)^{r_{0}-1}.$
@@
(4.17) $y_{\ell}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}.$
@@
$>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{12}c^{12}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{12r_{0}}d^{6r_{0}}y_{k}^{12r_{0}-3}.$
@@
(4.19) $\left(f_{k}f_{\ell}\right)^{8}\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(\frac{1.22\cdot 10^{7}}{\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$
@@
$\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}\left(1.22\cdot 10^{7}\right)^{2r_{0}-1}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}$
@@
$\frac{0.00078\left|N_{\alpha}\right|^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{d^{4}}>\left(\frac{y_{k}^{4}d^{2}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{8}}{59.2^{4}\left|N_{\alpha}\right|^{6}}\right)^{2r_{0}-1}.$